OpenStudy (anonymous):

Can someone help me , Solve this system of equations using the addition method. 5x 2y = 19 x - 2y = -1

OpenStudy (anonymous):

put 1 equation on top of the other one, in column 5x + 2y = 19 19x -2y = -1 and then add then in column to get 24x = 18 then solve for x and plug it back into either of the original equation to get y

OpenStudy (anonymous):

i got 3,2

OpenStudy (anonymous):

x=3 y=2

OpenStudy (anonymous):

is the first equation 5x + 2y = 19? You're missing something in the first equation

OpenStudy (anonymous):

yes thats the way its written on my quiz.

OpenStudy (anonymous):

5x2y? Doesn't make sense to me

OpenStudy (anonymous):

if it's 5x + 2y = 19 and x - 2y = -1, then the answer is (3, 2)